- Press Release
- August 15, 2022
ESMOS Status Meeting Minutes 4 Jun 2002
The ESMOS Status Meeting was conducted by K. Dolan at 1:00 PM June 4, 2002. The meeting was held in Room N202 in Building 32.
General Comments and Related Discussion:
The travel budget for Code 420 has been re-allocated; there was a $100K cut.
New promotion criteria have come out; additional training is required for managers. Details can be found on the Human Resources web site.
Award money for Code 400 personnel is being allocated; paperwork is due on Thursday.
New time and attendance guidelines are out.
There is a party for Diane Williams on June 18th.
R. J. Thompson is now the head of the EROS Data Center.
Alan will be in charge while Ken is attending a CALIPSO review on June 6th and 7th.
Discussion of ESMOS Top Ten Issues
#1 FY03 budget SOMO backing out of funding networks, institutional support and technology programs. Revamp programs. Cover increased security costs: No updates.
#2 Getting a contract in place to provide anomaly support from LMMS, SBRS, and L3COM: Revised statement of work for Lockheed is in place. It appears that purchase agreements will work for Santa Barbara.
#3 NISN/Code M line cost threat: No updates.
#4 End of Mission planning for ERBS and UARS: Trying to decommission ERBS by the end of the fiscal year. There was a meeting at NASA headquarters on the new iîªútute that is to be put in place with Capital College.
#5 Problem reprocessing ASTER data: There is a telecon with the Japanese tomorrow to discuss the issue; there will be a face-to-face meeting at JPL on June 24th and 25th.
#6 System vulnerabilities: No updates.
#7 Space in Building 3/14 for EDOS development facility: The space is pretty much ready for EDOS.
#8 Aqua Transition: Still waiting for the plan.
#9 DSMC transition NCC and WOTIS operations are moving to WSC during 2002: The ORR is scheduled for June 11th at White Sands; there will be a telecon connection at GCP. There was a NCC meeting on the transition last week. Manual tests will be scheduled for switching back and forth.
#10 USGS has not provided money to cover FY02 Landsat-7 network costs. NASA has not budgeted for this cost: Will be scheduling telecons at a higher level if the issue isnt resolved soon.
Comments by A. Johns and Related Discussion:
The ERBS SRR is scheduled for today.
CSOC is continuing to juggle their numbers.
Comments by B. Guit and Related Discussion:
Current activity is collision avoidance analysis for the upcoming Aqua burns.
Progress has been made on the Aqua transition plan. There is also progress on the TRW sustaining engineering support plan.
UARS and TOMS are going fine.
Comments by D. Lorenz and Related Discussion:
Been working on the ERBS SRR.
Comments by E. Corderman and Related Discussion:
Working on Aqua collision avoidance planning.
Started planning for Aura.
Comments by J. Wallace and Related Discussion:
Building 32 front entrance construction is starting on Thursday.
Comments by P. Johnson and Related Discussion:
The Aqua Build 11 content review is today; delivery is planned for early July.
Comments by L. Gayhart and Related Discussion:
This meeting will be held in Room W218 on July 9th.
Comments by E. Macie and Related Discussion:
The shuttle launch is on schedule for tomorrow.
Comments by R. Casey and Related Discussion:
What is the status of the AMSR-E SPS problem? Bill Guit stated that an anomaly meeting was held yesterday; issue is still under investigation.
Comments by C. Sigman and Related Discussion:
There is an EBnet upgrade design review tomorrow at 9:00 AM.
A Terra FDS security scan is planned for tomorrow after the burn is completed.